# Coordinated Inauthentic Behavior on TikTok: Challenges and Opportunities for Detection in a Video-First Ecosystem

### v1

@@ Results
Figure 4: Clustering of spectrogram vectors extracted from videos shared by coordinated accounts using identical hashtag sequences. The inset shows a t-SNE dimensionality reduction of the mel-spectrograms, highlighting distinct clusters corresponding to different AI-generated synthetic voices.
@@
Focusing on user activity in September, we applied the same approach to the corresponding Hashtag Sequence similarity network, yielding a dense network of 68 users with a graph density of 0.98—another clear example of collective similarity. In addition to repeatedly sharing the same hashtag sequences, users in this cluster also posted identical video content and descriptions. Notably, they employed a split-screen format, in which the left side of the video featured the actual content, typically political debate footage, while the right side showed unrelated visual material, such as chocolate-making (see Figure 5). While they resemble a duet, these videos are independently created and structured to imitate the format—possibly as an attempt to evade detection by obscuring the duplication of core content.
Figure 5: Examples of split-screen videos (left) and profiles (right) from coordinated accounts sharing identical hashtag sequences.
@@
Taken together, these findings, shared or inauthentic usernames, synchronized posting, identical content, evasion techniques, and synthetic voice replication, strongly suggest a centrally coordinated campaign designed to amplify specific political messaging through inauthentic means.
@@
All of these accounts included URLs in their video descriptions that redirected to the same website—shown in Figure 6 with candidate-specific details intentionally obscured for anonymity and privacy. Although the web domains varied slightly, they all ultimately pointed to the same landing page. Manual inspection revealed that the web page promotes a film associated with “CandidateY”. The landing page features a retro-styled embedded video player and a sign-up form requesting users’ names and email addresses in exchange for free access to the film. The page markets the film as available “for a limited time,” potentially encouraging urgency and engagement. This website was consistently linked across posts by accounts in the cluster.
@@
For September, pruning the network yielded a dense cluster of 33 users, many of whom also appeared in the August cluster (see the Appendix). These accounts exhibited nearly identical characteristics to those observed in August, but the number of such accounts had doubled. All usernames contained the string “CandidateY” and included the same URL in their video descriptions, linking to the external website shown in Figure 6. As before, the posts lacked hashtags and displayed synchronized posting patterns (see the Appendix).
@@
To identify coordinated communities, the network was pruned using a combination of edge filtering and node pruning (Cinus et al. 2025). First, edges with weights below the 99.5th percentile of similarity values were removed; then, node pruning was applied using a 95th percentile threshold—following the conservative criteria established in (Pacheco et al. 2021) and (Luceri et al. 2024), respectively. For August, the pruning process resulted in four isolated communities, as illustrated in Figure 7. Upon further investigation, several of these communities exhibited distinct forms of coordinated activity, which we detail as follows.
@@
Cluster 0: The largest cluster corresponds to the group of accounts promoting “CandidateY”, which we detected using the Co-Domain trace based on shared URLs in video descriptions. However, the time synchronization indicator analyzed here revealed an even broader set of accounts exhibiting similar characteristics—suggesting that not all users in this coordinated group relied on URLs to signal affiliation and promote content. All together, they consistently shared duplicate content within narrow time windows, indicating synchronized and likely centrally organized activity.
@@
Cluster 2: Accounts in this cluster appear to be part of a coordinated marketing campaign. The same individual appears across all videos, and usernames follow a consistent pattern—each beginning with “hd.since”. These users posted duplicated video content promoting both Republican- and Democrat-themed merchandise, suggesting an effort to monetize rather than promote a specific ideological stance.
@@
Focusing on user activity in September, we identified two prominent clusters. The first appears to represent a group engaged in coordinated campaigns promoting alleged donation efforts for families in Gaza. Each user in this cluster typically posts videos calling for support or donations, often framed as personal or urgent appeals. A notable tactic employed by this group involves starting each video with an unrelated pre-clip—such as trending TikTok content, humorous moments featuring popular creators, or visually satisfying footage. This strategy likely serves to capture viewer attention, extend watch time, and maximize engagement before revealing the actual fundraising message. In addition, the group frequently leveraged election-related keywords in video descriptions to potentially hijack TikTok’s recommendation algorithm and broaden content visibility. Videos across this cluster shared identical hashtags and descriptions, but typically directed users to different PayPal donation links, often placed in user bios. In some cases, identical PayPal links were reused across multiple accounts. Further signals of coordination included the reuse of profile pictures and a consistent username pattern, with most usernames ending in the suffix “family”. We also observed username changes and swapping among several accounts over time—a strategy previously linked to campaigns seeking to promote entities and run follow-back schemes (Pacheco et al. 2021; Mariconti et al. 2017). Collectively, these patterns suggest a likely case of inauthentic coordination and raise concerns about deceptive campaign practices on the platform.
@@
The second cluster comprises users whose usernames consistently begin with the name of “CandidateX”. These users regularly post TikTok videos that are highly similar in content and are synchronized in their timing, often appearing within minutes of each other. Notably, the videos shared by these users often consist of static images overlaid with political text, rather than original or dynamic audiovisual content. In many cases, the videos feature AI-generated voiceovers that narrate the on-screen text, contributing to a uniform and automated presentation style across accounts. This lack of media diversity, combined with the repeated use of a fixed set of hashtags, reinforces the hypothesis that these users are engaged in a coordinated campaign aimed at promoting “CandidateX”. Interestingly, 38.2% of users in this cluster also appear in the set of accounts identified through the hashtag sequence behavioral trace, indicating that a substantial proportion relied on identical or near-identical hashtag sequences to disseminate political messaging. While users identified through the hashtag sequence indicator tended to use a longer and more consistent hashtag order, users in this cluster employed a slightly shorter and reordered variant. Despite the variation in order, the core set of hashtags remained consistent, suggesting a possible effort to evade detection by subtly modifying posting patterns. Across this group, users demonstrated high similarity not only in hashtag usage but also in content structure, posting cadence, and strategic intent, all of which point to a deliberate, coordinated effort to promote support for “CandidateX”.
Challenges of CIB Detection with other behavioral signals and interaction patterns
@@
While the presented results offer substantial evidence of CIB across multiple dimensions, we also acknowledge that some indicators did not yield compelling evidence of coordination. Below, we describe the negative results uncovered during our analysis, highlighting the limitations and challenges associated with these alternative signals.
Co-Duet and Co-Stitch.
The Co-Duet and Co-Stitch networks were pruned using various combinations of eigenvector centrality and edge weight thresholds; however, we did not identify any suspicious clusters indicative of coordinated behavior. This is likely due to the fact that Duet and Stitch interactions are common practices on TikTok, often used to engage with popular content or express opinions on trending topics. In our dataset, users mainly performed Duet and Stitch to share their views on political discourse—a behavior more reflective of organic engagement than CIB. For both activities, our analysis revealed that users frequently employed these interaction modes to criticize political candidates or target prominent supporters. These videos often reuse segments from content originally posted by official accounts of public figures or media outlets, repurposing them to frame oppositional narratives. While this behavior did not exhibit the structural or temporal consistency typical of coordinated campaigns, it demonstrates how the Duet and Stitch features can be strategically leveraged by users across the political spectrum to deliver counter-messaging, even in the absence of organized coordination.
Co-Domain - Comments.
We examined the Co-Domain indicator by analyzing user comments to identify potential traces of CIB through URL sharing. The Co-Domain similarity network derived from user comments exhibited a notably sparse structure, consisting of multiple small, isolated components and a limited number of users, most of whom shared links to other TikTok videos rather than external websites. This approach did not reveal strong evidence of coordination, suggesting that TikTok users do not significantly engage in URL-based coordination via comments. This finding aligns with TikTok’s platform design, which emphasizes video and multimedia content over textual interaction.
Co-Reply.
We pruned the Co-Reply similarity network using various combinations of eigenvector centrality and edge-weight thresholds, consistently identifying a cluster of users who were highly engaged and frequently shared political content via replies. These users often responded to comments from opposing viewpoints, actively defending their own positions. A common trait, evident from their bios, was a strong and explicit alignment with a specific political side. Following their identification, we conducted a detailed analysis to investigate potential signs of coordination. Although the hashtag sequences used by these users were not identical, they often shared common hashtags, suggesting thematic alignment. However, there was no evidence of temporal synchronization in their posting activity. Further examination of user bios, profile pictures, and external links revealed no notable overlap or unifying features. Additionally, transcript similarity analysis did not indicate the presence of highly repetitive or closely aligned content. The most salient shared feature among this group was their political orientation, with the majority of posts aimed at criticizing the opposing political wing. While ideologically aligned, the absence of structural, temporal, or content-based similarities suggests that their behavior reflects organic partisanship rather than coordinated inauthentic activity.
Text Similarity - Comments.
We constructed similarity networks based on textual similarities in user comments and applied pruning using various combinations of eigenvector centrality and edge weight thresholds. Upon examining the resulting clusters, we observed that users often shared common keywords and conveyed aligned sentiments—for example, expressing support for or criticism of specific political parties. However, their behavior did not show any CIB-like activity. Rather, users within the same cluster tended to independently exhibit strong and polarized opinions. An additional insight is that this approach, focused solely on textual similarity in comments, does not incorporate the underlying video content. As a result, comments made on the same video, especially those posted by prominent news channels, tended to exhibit higher textual similarity, likely due to shared contextual triggers. We also noted a high degree of overlap in the videos commented on by users in certain clusters, suggesting clustering was driven more by common exposure to news content than by coordinated engagement.
Text Similarity - Transcripts.
We applied the same strategy used for comment similarity to transcript similarity, i.e., the spoken content in the videos. The resulting graph remained very dense, even after applying stringent pruning criteria. This density can be attributed to common practices on TikTok, where users frequently incorporate snippets from political speeches, news footage, and other widely circulated media. As a result, even with high similarity thresholds, the number of accounts connected within clusters remained substantially large, limiting the method’s effectiveness for isolating CIB clusters. Manual inspection revealed that users frequently reused the same media clips, often sourced from news outlets, political campaigns, or Duet and Stitch interactions with political commentators. Notable examples include clips from Donald Trump’s campaign rallies and footage of Beyoncé at a Kamala Harris event. While these behaviors reflect strong thematic convergence, they appear to be the product of platform norms and organic engagement, rather than evidence of structured or deceptive coordination.
@@ Discussion
This study introduces the first large-scale computational framework for detecting coordinated inauthentic behavior (CIB) on TikTok—a platform that has been largely unexplored in CIB research despite its growing influence in political discourse and social issues. Our work adapts and extends network-based coordination detection methods originally developed for text-centric platforms like Twitter/$\mathbb{X}$ to TikTok’s unique multimodal and interaction environment. By leveraging behavioral traces such as hashtag sequences, posting synchronicity, shared external domains, and content reuse, we constructed and pruned user similarity networks to uncover clusters of potentially coordinated users.
Our findings demonstrate that traditional behavioral signals like hashtag sequence overlap, external web pages amplification, and synchronized posting remain effective indicators of coordination on TikTok. While studying a dataset related to the 2024 U.S. Presidential Election, we uncovered numerous instances of inauthentic activity, including semi-automated political messaging campaigns, orchestrated AI-assisted content creation, and artificial amplification through shared web domains. These signals, when converging across multiple traces, strongly suggest centrally organized activity aimed at exploiting TikTok’s recommendation algorithm to boost political narratives or engagement-driven campaigns. However, only a small fraction of the CIB accounts we identified appear to have been removed from the platform. This highlights a potential gap between the detection of coordinated manipulation and TikTok’s enforcement actions, underscoring the opportunity for more proactive and robust moderation strategies to detect and disrupt such activity at scale. This disconnect between detection and enforcement is particularly salient in light of TikTok’s unique content environment, which presents both challenges and opportunities for coordination detection. While platform-native features like Duet, Stitch, and the reuse of political video segments tend to blur the line between organic engagement and coordination, our framework uncovered more subtle and emerging signals of inauthentic behavior on TikTok, including watermark reuse, synthetic voice patterns, and manufactured split-screen videos. Together, these findings highlight the importance of developing computational models that reflect the evolving tactics and increasing sophistication of AI-enabled inauthentic coordination on video-first platforms.
@@
This study has several important limitations. First, while our method effectively identified coordination using several behavioral signals, others did not yield clusters indicative of inauthentic behavior. Future work is needed to assess whether this outcome reflects limitations in our current methods for capturing these signals, a lower prevalence of their deceptive use in CIB on TikTok, or characteristics specific to the dataset and discussion context analyzed. Second, we found that no single pruning or filtering strategy was consistently effective across all indicators, suggesting that the heterogeneity of user behavior on TikTok requires adaptive, trace-specific tuning. Third, our conservative approach prioritizes precision over recall, providing a lower bound on the prevalence of potential CIB-like activity. While this minimizes false positives, it may also overlook weaker or emerging coordination signals. This trade-off is evident in our results, where newly active campaigns in September obscure the patterns identified in August. Lowering thresholds would improve continuity but may introduce noisier clusters with less compelling evidence of CIB. Finally, while similar user clusters emerged across different indicators, they were not fully overlapping, suggesting that the signals capture complementary aspects of coordination. Future work will explore fusion strategies that integrate multiple indicators.
@@
We recognize both the positive and negative societal implications of CIB detection efforts. On the one hand, this research contributes to safeguarding the integrity of online discourse and offers computational tools to identify covert manipulation campaigns, potentially supporting democratic resilience. On the other hand, there are risks of misapplication, including the inadvertent labeling of organic, legitimate activism as inauthentic, or the weaponization of detection frameworks to suppress dissent. To mitigate these risks, we designed our framework to prioritize transparency and explainability, and we emphasize cautious interpretation of coordination signals, particularly in contexts involving political or activist speech.

### v2

@@ Results
Figure 4: Clustering of spectrogram vectors extracted from videos shared by coordinated accounts using identical hashtag sequences. The inset shows a t-SNE dimensionality reduction of the spectrograms, highlighting distinct clusters corresponding to different AI-generated synthetic voices.
@@
Focusing on user activity in September, we applied the same approach to the corresponding Hashtag Sequence similarity network, yielding a dense network of 68 users (a completely different set from the one found in August) with a graph density of 0.98—another clear example of collective similarity. In addition to repeatedly sharing the same hashtag sequences, users in this cluster also posted identical video content and descriptions. Notably, they employed a split-screen format, in which the left side of the video featured the actual content, typically political debate footage, while the right side showed unrelated visual material, such as chocolate-making (see Figure 5). While they resemble a duet, these videos are independently created and structured to imitate the format, possibly as an attempt to evade detection by obscuring the duplication of core content.
Figure 5: Examples of split-screen videos and profiles from coordinated accounts sharing identical hashtag sequences.
@@
In October, we identified a coordinated network of 19 users, now all removed or suspended from TikTok, who engaged in strikingly synchronized behavior. These accounts posted 157 videos featuring identical hashtag sequences and nearly identical descriptions that included “CandidateX” name and a numeric identifier that varied systematically between posts, suggesting the use of a templated posting strategy. Activity peaked on October 17th and 18th, with many videos published within tight time windows. Taken together, the combination of identical hashtags, near-duplicate descriptions, and synchronized posting provides strong evidence of coordinated inauthentic behavior.
@@
All of these accounts included URLs in their video descriptions that redirected to the same website, shown in Figure 6 with candidate-specific details intentionally obscured for anonymity and privacy. Although the web domains varied slightly, they all ultimately pointed to the same landing page. Manual inspection revealed that the web page promotes a film associated with “CandidateY”. The landing page features a retro-styled embedded video player and a sign-up form requesting users’ names and email addresses in exchange for free access to the film. The page markets the film as available “for a limited time,” potentially encouraging urgency and engagement. This website was consistently linked across posts by accounts in the cluster.
@@
For September, pruning the network yielded a dense cluster of 33 users, many of whom also appeared in the August cluster (see the Appendix). These accounts exhibited nearly identical characteristics to those observed in August, but the number of such accounts had doubled. All usernames contained the string “CandidateY” and included the same URL in their video descriptions, linking to the external website shown in Figure 6. As before, the posts lacked hashtags and displayed synchronized posting patterns (see the Appendix). In October, we found no evidence of coordinated behavior using the co-domain behavioral trace, suggesting that this network either shifted its tactics or was dismantled. Notably, as of today, all accounts previously involved in promoting “CandidateY” have been removed or suspended from TikTok, potentially indicating moderation interventions.
@@
To identify coordinated clusters with the synchronous posting behavioral trace, the network was pruned using a combination of edge filtering and node pruning (Cinus et al. 2025). First, edges with weights below the 99.5th percentile of similarity values were removed; then, node pruning was applied using a 95th percentile threshold, following the conservative criteria established in (Pacheco et al. 2021) and (Luceri et al. 2024), respectively. For August, the pruning process resulted in four isolated clusters, as illustrated in Figure 7. Upon further investigation, all these clusters exhibited distinct forms of coordinated activity, which we detail as follows.
@@
Cluster 0: The largest cluster corresponds to the group of accounts promoting “CandidateY”, which we detected using the Co-Domain trace based on shared URLs in video descriptions. However, the time synchronization indicator analyzed here revealed an even broader set of accounts exhibiting similar characteristics, suggesting that not all users in this coordinated group relied on URLs to signal affiliation and promote content. All together, they consistently shared duplicate content within narrow time windows, indicating synchronized and likely centrally organized activity.
@@
Cluster 2: Accounts in this cluster appear to be part of a coordinated marketing campaign. The same individual appears across all videos, and usernames follow a consistent pattern, each beginning with “hd.since”. These users posted duplicated video content promoting both Republican- and Democrat-themed merchandise, suggesting an effort to monetize rather than promote a specific ideological stance.
@@
Focusing on user activity in September, we identified two prominent clusters. The first appears to represent a group engaged in coordinated campaigns promoting alleged donation efforts for families in Gaza. Each user in this cluster typically posts videos calling for support or donations, often framed as personal or urgent appeals. A notable tactic employed by this group involves starting each video with an unrelated pre-clip, such as trending TikTok content, humorous moments featuring popular creators, or visually satisfying footage. This strategy likely serves to capture viewer attention, extend watch time, and maximize engagement before revealing the actual fundraising message. In addition, the group frequently leveraged election-related keywords in video descriptions to potentially hijack TikTok’s recommendation algorithm and broaden content visibility. Videos across this cluster shared identical hashtags and descriptions, but typically directed users to different PayPal donation links, often placed in user bios. In some cases, identical PayPal links were reused across multiple accounts. Further signals of coordination included the reuse of profile pictures and a consistent username pattern, with most usernames ending in the suffix “family”. We also observed username changes and swapping among several accounts over time—a strategy previously linked to campaigns seeking to promote entities and run follow-back schemes (Pacheco et al. 2021; Mariconti et al. 2017). Collectively, these patterns suggest a likely case of inauthentic coordination and raise concerns about deceptive campaign practices on the platform.
@@
The second cluster comprises users whose usernames consistently begin with the name of “CandidateX”. These users regularly post TikTok videos that are highly similar in content and are synchronized in their timing, often appearing within minutes of each other. Notably, the videos shared by these users often consist of static images overlaid with political text, rather than original or dynamic audiovisual content. In many cases, the videos feature AI-generated voiceovers that narrate the on-screen text, contributing to a uniform and automated presentation style across accounts. This lack of media diversity, combined with the repeated use of a fixed set of hashtags, reinforces the hypothesis that these users are engaged in a coordinated campaign aimed at promoting “CandidateX”. Interestingly, some of the users in this cluster also appear in the set of accounts identified through the hashtag sequence behavioral trace. While users identified through the hashtag sequence indicator tended to use a longer and more consistent hashtag order, users in this cluster employed a slightly shorter and reordered variant. Despite the variation in order, the core set of hashtags remained consistent, suggesting a possible effort to evade detection by subtly modifying posting patterns. Across this group, users demonstrated high similarity not only in hashtag usage but also in content structure, posting cadence, and strategic intent, all of which point to a deliberate, coordinated effort to promote support for “CandidateX”.
Figure 8: UMAP projection of audio embeddings: Each point corresponds to an individual audio clip, sourced either from verified, authentic recordings or from TikTok videos shared by coordinated accounts.
CIB Detection via Speech Similarity
For the month of August, the Speech Similarity network includes 102 accounts, all of which feature the string “CandidateY” in their usernames. This naming convention aligns with accounts previously identified through co-domain and synchronous posting traces. Although this coordinated network likely belongs to the broader campaign promoting “CandidateY”, its accounts and tactics, centered on repeatedly sharing videos with identical speech segments, differ significantly from those identified in previous analyses.
The Speech Similarity network for September revealed a cluster of 42 accounts. These accounts posted videos with identical audio content at the exact same time, though each video featured distinct visuals. The content often promoted motivational messaging associated with “CandidateX”, and account usernames followed consistent themes, combining “CandidateX” name with words like motivation, mindset, lesson, or advices. These usernames and the resulting network structure are shown in the Appendix. Manual inspection reveals a recurring video format: each post begins with a photo of “CandidateX”, followed by unrelated footage, such as scenic landscapes, inspirational imagery, or abstract visuals (examples in the Appendix). Importantly, “CandidateX” appears only within the first five seconds of each video, suggesting a strategy of attention capture followed by narrative redirection, a tactic frequently employed in influence campaigns seeking to optimize user retention.
In addition to visual consistency, a key commonality across these videos is the use of speech segments that resemble “CandidateX” voice. To assess whether these were authentic or synthetic, we conducted an audio-based forensic analysis using the ECAPA-TDNN model (Desplanques, Thienpondt, and Demuynck 2020), which is designed to capture subtle vocal cues and has been validated for detecting AI-generated or manipulated speech. We compared 426 suspect audio clips from TikTok videos shared by the coordinated network against 3,268 verified recordings of “CandidateX” from the In the Wild dataset (Müller et al. 2022). All files were standardized, resampled to 16 kHz, and encoded via a pretrained ECAPA model to obtain audio embeddings. The resulting embeddings, visualized in Figure 8 via UMAP dimensionality reduction, reveal a clear separation between verified and suspect audio samples (Fisher’s ratio $\approx$ 5.77). While a few TikTok clips may plausibly feature genuine speech, the overall distribution provides strong evidence that a substantial portion of the suspect audios were artificially synthesized. Manual annotations confirmed these results (see the Appendix for more details). This suggests a deliberate strategy to craft persuasive or emotionally resonant content, highlighting the evolving tactics of inauthentic campaigns that exploit the affordances of short-form video and synthetic media.
The coordinated network resurfaced in October, alongside a separate spam campaign that hijacked election-related hashtags to promote a gaming application.
Figure 9: Probability density function of posting time differences and daily activity patterns of coordinated (video similarity network) vs. organic accounts
CIB Detection via Video Similarity
Through our video similarity framework, we uncovered a network of 67 coordinated accounts active in August. These accounts uniformly included the string “CandidateY” in their usernames, consistent with those identified through other behavioral traces such as speech similarity, synchronous posting, and co-domain. To quantify the overlap across these sets of coordinated accounts, we compute the Normalized Mutual Information (NMI) scores among the four networks. The consistently low values (see the Appendix for details) confirm that, even within a single campaign, distinct sets of accounts engage in different coordination strategies. This aligns with prior research showing that influence campaigns often deploy multiple, minimally overlapping subnetworks to evade detection (Luceri et al. 2024; Ng and Carley 2022).
Figure 9 visualizes the temporal signature of the coordinated activity of accounts in the video similarity network, comparing it against a baseline of organic (non-coordinated) accounts. The main panel displays the probability density of time gaps between their posts: coordinated accounts exhibit a sharp peak at 0 hours followed by a steep decline, while organic accounts show a broader, more gradual distribution. The inset further highlights the hourly activity patterns, revealing that coordinated accounts post almost exclusively between 6pm and 8pm UTC (2pm–5pm ET), in contrast to the more evenly distributed posting times of organic users, strengthening the evidence of inauthentic, synchronized amplification. The video similarity network for September and October continued to surface the same accounts uncovered in August. All the accounts in these clusters are no longer active on TikTok, possibly following moderation actions.
Challenges of CIB Detection on TikTok
TikTok’s Research API poses structural limitations, such as rate limits, incomplete metadata returns, and inconsistent quota fulfillment, that can hinder full behavioral coverage (Corso, Pierri, and De Francisci Morales 2024; Entrena-Serrano et al. 2025). To account for this, we simulated 5% and 10% data loss and found our detection methods remained robust, recovering nearly all coordinated accounts (see the Appendix for more details). Mitigating these limitations at scale requires hybrid pipelines that combine the API with compliant third-party tools, as well as efficient sampling strategies and distributed crawling infrastructure.
@@
While our framework uncovered several robust indicators of CIB, some behavioral traces did not reveal compelling evidence of coordination. These negative results are important for understanding the limitations and challenges associated with these alternative signals.
The analysis of Duet and Stitch interactions revealed no suspicious clusters indicative of coordinated behavior, even after applying network pruning techniques based on eigenvector centrality and edge weight. This is likely because these features are organically used on TikTok to engage with popular content or express opinions on trending topics. In our dataset, users mainly performed Duet and Stitch to share their views on political discourse—a behavior more reflective of organic engagement than CIB.
In examining co-domain signals through user comments, we found that the resulting network was sparse and fragmented. Most of the shared links pointed to other TikTok videos rather than external domains. This aligns with TikTok’s platform design, which de-emphasizes URL sharing and textual interaction. As a result, URL-based coordination through comments appears to be rare and did not provide useful insights into inauthentic behavior.
The co-reply network yielded a cluster of highly active users engaged in political conversations, often replying to opposing viewpoints and defending specific candidates. While these users shared common hashtags and displayed strong ideological alignment, the lack of temporal synchronization, structural cohesion, or repetitive messaging suggests that their behavior was driven by organic partisanship rather than orchestrated coordination.
No substantial evidence of coordination emerged from textual similarity patterns in user comments. Although users often expressed polarized opinions and used similar language, these similarities were more attributable to commenting on the same trending videos than to coordinated messaging. We also noted a high degree of overlap in the videos commented on by users in certain clusters, suggesting clustering was driven more by common exposure than by coordinated engagement.
@@ Discussion
[…]
By leveraging diverse behavioral signals, including hashtag sequences, synchronized posting, shared web domains, and multimodal content reuse, we construct similarity networks that expose clusters of potentially coordinated users. Applied to a dataset centered on the 2024 U.S. Presidential Election, our approach reveals multiple instances of inauthentic activity, ranging from semi-automated political messaging campaigns to synthetic media usage and orchestrated content duplication. These signals, especially when converging across traces, indicate structured efforts to exploit TikTok’s algorithmic affordances to amplify political narratives.
Despite uncovering numerous examples of likely CIB, we find that many accounts involved in such behavior remain active on the platform. This highlights a potential gap between the detection of coordinated manipulation and enforcement actions, underscoring the need for proactive and scalable interventions tailored to TikTok’s ecosystem.
Key Takeaways.
Our analysis of TikTok-specific interaction patterns yields several actionable insights:
TikTok’s ecosystem fosters new and traditional modalities of coordination: We observed both conventional strategies, like co-posting of URLs or copy-paste textual duplication, and TikTok-specific behaviors, such as manufactured split-screen videos and visuals featuring synthetic voiceovers. The cross-modal reuse of content (e.g., the same audio with different videos, and vice versa) suggests that TikTok’s algorithmic affordances are being actively gamed to amplify messaging while avoiding detection. Coordinated behaviors are increasingly adapted to the interaction norms and remix practices that define the platform’s user culture.
Multimodal coordination requires multimodal detection: Many coordinated behaviors on TikTok—such as synchronized posting of near-identical videos or the use of synthetic voiceovers—would likely evade traditional, text-centric detection models. Unlike platforms where coordination often involves direct duplication of text or links, TikTok campaigns frequently rely on templated strategies and multimodal reuse: swapping visuals while retaining the same audio track, or employing split-screen formats where political content appears alongside unrelated, visually engaging footage. Detecting such tactics necessitates incorporating both visual and audio-based similarity metrics.
The growing role of Generative AI in CIB: One of our key findings is the emergence of AI-generated voiceovers and synthetic speech used to simulate political figures and deliver persuasive messaging with minimal human labor. This trend marks a pivotal shift in coordination tactics: inauthentic campaigns are increasingly turning to generative AI to craft emotionally resonant, high-reach media at scale, sidestepping traditional, text-centric influence strategies.
CIB detection on TikTok is feasible, but coordination signals are highly fragmented: Our framework successfully uncovered multiple CIB clusters linked to influence campaigns, yet each signal tended to expose largely disjoint sets of accounts. Even within the same campaign, different groups employed distinct tactics. This underscores that TikTok CIB does not coalesce around a single behavior, but instead relies on complementary strategies, requiring detection pipelines that integrate multiple weak signals.
Not all interaction features are useful: Duets, stitches, and video-based replies did not produce strong signals of coordination. These are often used for reactive or oppositional content creation, not organized manipulation.
@@
This study has several important limitations. First, while our method effectively identified coordination using several behavioral signals, others did not yield clusters indicative of inauthentic behavior. Future work is needed to assess whether this outcome reflects limitations in our current methods for capturing these signals, a lower prevalence of their deceptive use in CIB on TikTok, or characteristics specific to the dataset and discussion context analyzed. Second, we found that no single pruning or filtering strategy was consistently effective across all indicators, suggesting that the heterogeneity of user behavior on TikTok requires adaptive, trace-specific tuning. Third, our conservative approach prioritizes precision over recall, providing a lower bound on the prevalence of potential CIB-like activity. While this minimizes false positives, it may also overlook weaker or emerging coordination signals. Finally, while similar user clusters emerged across different indicators, they were not fully overlapping, suggesting that the signals capture complementary aspects of coordination. Future work will explore fusion strategies that integrate multiple indicators.
@@
We recognize both the positive and negative societal implications of CIB detection efforts. On the one hand, this research contributes to safeguarding the integrity of online discourse and offers computational tools to identify covert manipulation campaigns, potentially supporting democratic resilience. On the other hand, there are risks of misapplication, including the inadvertent labeling of organic, legitimate activism as inauthentic, or the weaponization of detection frameworks to suppress dissent. To mitigate these risks, we designed our framework to prioritize transparency and explainability, and we emphasize cautious interpretation of coordination signals.
@@
Impact of Data Loss on Detection Strategies: A Robustness Assessment
We assessed the robustness of our detection strategies under simulated data loss scenarios, randomly removing 5% and 10% of video data. Our method proved highly resilient, consistently retaining nearly all coordinated accounts under both data loss conditions.
  • Hashtag Sequence: Retained 98.5% of coordinated accounts under both 5% and 10% data loss.
  • Co-Domain: Retained 98.5% of coordinated accounts under 5% loss and 97.1% under 10% loss.
  • Speech Similarity: Retained 98.0% of coordinated accounts under 5% loss and 91.2% under 10% loss.
  • Video Similarity: Retained 97.1% of coordinated accounts under 5% loss and 92.6% under 10% loss.
Manual Annotations Inter-Agreement
Manual inspections were conducted by two annotators across several categories to validate the reliability of our findings. For annotations assessing whether usernames in distinct clusters contained references to either CandidateX or CandidateY, we observed perfect agreement (Cohen’s Kappa = 1.0). The same level of inter-annotator agreement was achieved when labeling: (i) synthetic vs. authentic voiceovers in CandidateX videos identified via the hashtag sequence indicator, (ii) Gaza-related accounts exhibiting handle rotation detected through synchronous posting, and (iii) manufactured split-screen versus duet videos.
Annotations of synthetic voices mimicking CandidateX (authentic vs. fake) yielded a Cohen’s Kappa score of 0.756, indicating substantial agreement. However, the strong class imbalance (most labels falling into the same category, i.e., fake) suppresses the Kappa value. The actual observed agreement was 421 out of 426 samples ($\sim$0.99), confirming that disagreement was extremely rare.
@@
Figure 15: Coordinated cluster of accounts spreading “CandidateX” messaging. Usernames across the network display uniform prefixes and motivational themes.
@@
Figure 17: Cluster of coordinated accounts sharing near-identical video transcripts (left panel) and visuals (right panel). Usernames within both clusters show strong similarity, all beginning with a prefix linked to a U.S. presidential candidate, denoted here as “CandidateY” to pre- serve anonymity and privacy. To reduce visual clutter, we display only the usernames of accounts that posted more than ten videos. Despite sharing similar usernames, the two networks exhibit minimal overlap (NMI = 0.1).
Keywords and Hashtags used in the Data Collection
  • election2024, Election2024
  • US Elections, USElections, us elections, uselections
  • 2024Elections, 2024 Elections, 2024 elections, 2024elections, 2024election
  • 2024PresidentialElections, 2024 Presidential Elections, 2024presidentialelections, 2024 presidential elections
  • saveamerica2024
  • presidentbiden, Biden, biden, bidenharris, JoeBiden, Joe Biden, joebiden, joe biden, joseph biden, Joseph Biden
  • Biden2024, biden2024, bidenharris2024
  • Donald Trump, donald trump, DonaldTrump, donaldtrump, donaldtrump2024
  • Trump2024, trump2024, trumpsupporters, trumptrain
  • republicansoftiktok, conservative
  • MAGA, maga, makeamericagreatagain, ultramaga, KAG
  • Republican, presidenttrump, trumpismypresident, letsgobrandon
  • GOP, CPAC
  • NikkiHaley, nikkihaley
  • DeSantis, RonDeSantis, desantis, rondesantis
  • RNC
  • democratsoftiktok, democratsarehot, thedemocrats
  • voteblue2024, vote blue
  • DNC, dnc
  • kamalaharris, kamala harris
  • mariannewilliamson, williamson2024
  • deanphillips, phillips2024
  • democratic party, Democratic party
  • republican party, Republican party
  • Third party, third party
  • Green party, green party
  • Independent party, independent party
  • RFKJr, RFK Jr., RFK Jr, rfkjr, rfkjr., rfk, Robert F. Kennedy Jr., robert f. kennedy Jr.
  • jill stein, jillstein, Jill Stein, JillStein
  • CornellWest, cornellwest
Figure 18: NMI score for CandidateY accounts detected across multiple strategies
Figure 19: Examples of TikTok videos from a coordinated network employing synthetic “CandidateX” voiceovers to deliver motivational messages.